OpenStudy (anonymous):

How do you solve \[2 \cos x \sin x - \sin x=0\] for \[ 0 \le x <2\pi\] I solved and got \[\frac{ \pi }{ 3 }\ and \frac{ 5\pi }{ 3 }\] My teacher said I should also include \[0\] and \[\pi\] Can anyone explain to me why?

OpenStudy (zehanz):

If you factor out the sinx you get:\[2\cos x \sin x-\sin x=0 \Leftrightarrow \sin x(2\cos x-1)=0\]Now, a product is zero if one or more factors of it are zero, so:\[\sin x=0 \vee 2\cos x -1 =0\] sin x = 0 has 0, pi and 2pi as solutions.

OpenStudy (zehanz):

2pi not included, because x < 2pi

OpenStudy (zehanz):

You probably divided by sin x, thereby letting two more solutions magically disappear!

OpenStudy (anonymous):

That is exactly what I did. Thanks. How would I know not to do that?

OpenStudy (zehanz):

Dividing by something unknown is not a good idea: sin x can be anything between -1 and 1. Mostly it is not equal to zero, but if it is, you can't divide by it! And when it is 0 ther you have a solution: Just remember if you've got an equation of the form ab=0: \[a \cdot b=0 \Leftrightarrow a=0 \vee b=0\]

OpenStudy (anonymous):

Awesome explanation - if I could give you a hundred medals, I would!

OpenStudy (zehanz):

Hear, hear...